# Rigidity and nonexistence of CMC hypersurfaces in 5-manifolds

### v2

@@ 4.2. Diameter bound
$|\nabla h|\leqslant\varepsilon_{0}\left(1+\frac{h^{2}}{2\varepsilon_{0}}\right).$
We now consider the functional $\mathcal{A}_{k}$ on $M^{4}$, where the weight function $u$ satisfies
$\begin{split}-\Delta_{M}u&=\left(|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)\right)u~{}~{}\mbox{in}~{}~{}\mathring{M},\\ u&=1~{}~{}\mbox{on}~{}~{}\partial M;\end{split}$
see [Wu23, Hong24] for more details on the construction of $u$. We assume that $\Omega$ is the minimizer of $\mathcal{A}_{k}$ and $\Sigma$ is a connected component of $\partial\Omega\backslash\partial M$. $\nu$ is the unit normal vector to $\Sigma$.
To eliminate the term $\psi^{2}hu^{k-1}\nabla_{\nu}u$ in the second variation of $\Sigma$, note that by Lemma 4.3
$u^{k}H_{\Sigma}^{2}\psi^{2}=k^{2}u^{k-2}\left(\nabla_{\nu}u\right)^{2}\psi^{2}+u^{k}h^{2}\psi^{2}-2ku^{k-1}h\psi^{2}\nabla_{\nu}u.$
Since $k\leqslant 2$, plugging into the stability inequality yields
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\geqslant&\int_{\Sigma}\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-\frac{1}{2}H_{\Sigma}^{2}\right)u^{k}\psi^{2}\\ &\int_{\Sigma}-ku^{k-1}\psi^{2}\Delta_{M}u+\psi^{2}u^{k}\nabla_{\nu}h+\frac{1}{2}u^{k}h^{2}\psi^{2}.\end{split}$
@@
(4.1) $-ku^{-1}\Delta_{M}u+\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-\frac{1}{2}H_{\Sigma}^{2}\right).$
We select a local orthonormal frame $\{e_{i}\}_{i=1}^{3}$ on $\Sigma^{3}$ and extend it on $X^{5}$ by assigning $\nu=e_{4}$ and $\eta=e_{5}$. By Gauss equation, we have
$\begin{split}\operatorname{Ric}^{\Sigma}(e_{1},e_{1})&=\sum_{i=1}^{3}\operatorname{Rm}^{\Sigma}_{i11i}=\sum_{i=1}^{3}\left(\operatorname{Rm}^{M}_{i11i}+A^{\Sigma}_{ii}A^{\Sigma}_{11}-\left(A^{\Sigma}_{i1}\right)^{2}\right)\\ &=\sum_{i=1}^{3}\operatorname{Rm}^{M}_{i11i}+H_{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{3}\left(A^{\Sigma}_{i1}\right)^{2}\\ &=\sum_{i=1}^{3}\operatorname{Rm}^{M}_{i11i}+\sum_{i=1}^{4}\operatorname{Rm}^{M}_{i44i}-\operatorname{Ric}^{M}(\nu,\nu)+H_{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{3}\left(A^{\Sigma}_{i1}\right)^{2}.\end{split}$
@@
$\begin{split}k\operatorname{Ric}^{X}&(\eta,\eta)+\sum_{i=1}^{3}\operatorname{Rm}^{M}_{i11i}+\sum_{i=1}^{4}\operatorname{Rm}^{M}_{i44i}\\ =&k\sum_{i=1}^{5}\operatorname{Rm}^{X}_{i55i}+\sum_{i=1}^{3}\left(\operatorname{Rm}^{X}_{i11i}+A^{M}_{ii}A^{M}_{11}-\left(A^{M}_{i1}\right)^{2}\right)\\ &+\sum_{i=1}^{4}\left(\operatorname{Rm}^{X}_{i44i}+A^{M}_{ii}A^{M}_{44}-\left(A^{M}_{i4}\right)^{2}\right)\\ =&k\sum_{i=1}^{5}\operatorname{Rm}^{X}_{i55i}+\sum_{i=1}^{3}\operatorname{Rm}^{X}_{i11i}+\sum_{i=1}^{4}\operatorname{Rm}^{X}_{i44i}\\ &+(H_{M}-A^{M}_{44})A^{M}_{11}-\sum_{i=1}^{4}\left(A^{M}_{i1}\right)^{2}+(A^{M}_{41})^{2}+H_{M}A^{M}_{44}-\sum_{i=1}^{4}\left(A^{M}_{i4}\right)^{2}\\ \geqslant&\lambda_{k}+(H_{M}-A^{M}_{44})A^{M}_{11}-\sum_{i=1}^{4}\left(A^{M}_{i1}\right)^{2}+(A^{M}_{41})^{2}+H_{M}A^{M}_{44}-\sum_{i=1}^{4}\left(A^{M}_{i4}\right)^{2}.\end{split}$
Besides, by Proposition 4.5, all terms involving $A_{M}$ in (4.1) can be estimated as follows:
$k|A_{M}|^{2}-A^{M}_{11}A^{M}_{44}-\sum_{i=1}^{4}\left(A^{M}_{i4}\right)^{2}-\sum_{i=2}^{4}\left(A^{M}_{i4}\right)^{2}+H_{M}(A^{M}_{11}+A^{M}_{44})\geq\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}.$
It is not hard to check that
$|A_{\Sigma}|^{2}+H_{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{3}\left(A^{\Sigma}_{i1}\right)^{2}\geqslant\frac{1}{2}H_{\Sigma}^{2}.$
In conclusion, we have the following inequality
@@
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\geqslant&\int_{\Sigma}\left(\varepsilon_{0}-\operatorname{Ric}^{\Sigma}(e_{1},e_{1})\right)u^{k}\psi^{2}\\ &+\int_{\Sigma}\left(\varepsilon_{0}+\frac{1}{2}h^{2}+\nabla_{\nu}h\right)u^{k}\psi^{2}\end{split}$
We take $\psi=u^{-\frac{k}{2}}\phi$. Direct computation shows
$\nabla^{\Sigma}\psi=u^{-\frac{k}{2}}\nabla^{\Sigma}\phi-\frac{k}{2}u^{-\frac{k}{2}-1}\phi\nabla^{\Sigma}u,$
and
$u^{k}|\nabla^{\Sigma}\psi|^{2}=|\nabla^{\Sigma}\phi|^{2}+\frac{k^{2}}{4}|\nabla^{\Sigma}\log u|^{2}\phi^{2}-k\phi\langle\nabla^{\Sigma}\log u,\nabla^{\Sigma}\phi\rangle.$
Integration by parts and Cauchy–Schwarz inequality yield
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}&-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\\ &\leqslant\left(k\left(\frac{\varepsilon}{2}-1\right)+\frac{k^{2}}{4}\right)\int_{\Sigma}u^{-2}|\nabla^{\Sigma}u|^{2}\phi^{2}+\left(1+\frac{1}{2\varepsilon}\right)\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\\ &\leqslant\frac{4}{4-k}\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\end{split}$
@@
by choosing $\varepsilon=2-\frac{k}{2}$. Choosing $e_{1}$ so that the minimum Ricci curvature scalar $\lambda_{\operatorname{Ric}}(\Sigma)=\operatorname{Ric}_{\Sigma}(e_{1},e_{1})$, we obtain that
@@
$\begin{split}\frac{4}{4-k}\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}&\geqslant\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\\ &\geqslant\int_{\Sigma}\left(-\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)\phi^{2}.\end{split}$
@@
Therefore, there is a smooth function $V\in C^{\infty}(\Sigma)$ so that
$V\geqslant\frac{4-k}{4}\left(-\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)$
and
$\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\geqslant\int_{\Sigma}V\phi^{2},$
for all $\phi\in C^{0,1}(\Sigma)$. Following the argument in [Chodosh-Li-Minter-Stryker][Theorem 5.1], the first eigenfunction of the operator $-\Delta_{\Sigma}-V$ is a smooth positive function $\theta$ that satisfies
@@
$-\Delta_{\Sigma}\theta\geqslant\frac{4-k}{4}\left(-\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)\theta.$
Let $\alpha=\frac{4}{4-k}$. Hence,
$\operatorname{Ric}^{(\theta,\alpha)}:=\operatorname{Ric}^{\Sigma}-\alpha^{-1}\left(\theta^{-1}\Delta_{\Sigma}\theta\right)g_{\Sigma}\geqslant\varepsilon_{0}g_{\Sigma}.$
Since $\alpha\leqslant 2$, we have by [shenyingyerugang][Corollary 1]
$\operatorname{diam}(\Sigma,g_{\Sigma})\leqslant\frac{2\pi}{\sqrt{2\varepsilon_{0}}}.$
Lemma 4.7.
Let $(X^{5},g)$ be a complete $5$-dimensional Riemannian manifold. Let $M^{4}\hookrightarrow(X^{5},g)$ be a (strongly) stable two-sided CMC hypersurface of the mean curvature $H_{M}$ with nonsmooth boundary $\partial M=\partial_{0}M\cup\partial_{1}M$, and $\partial_{1}M$ intersects $\partial_{0}M$ at angle $(0,\pi/8)$. Suppose
$\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\varepsilon_{0}$
for a positive constant $\varepsilon_{0}$ and some $k\in[\frac{3}{2},2]$, where $\lambda_{k}$ is the infimum of the $k$-triRicci curvature. There are universal constants $L>0$ and $c>0$ such that if there is a $p\in M$ with $d_{M}(p,\partial M)>\frac{L}{2}$, then in $M$ there is a relatively open set $\Omega\subset B^{M}_{L/2}(\partial_{1}M)$ and a smooth $3$-dimensional Riemannian manifold $\Sigma^{3}$ with boundary (possibly disconnected) such that $\Sigma=\partial\Omega\cap M^{\circ}$ and each component of $\Sigma$ has diameter at most $c$.
The proof of this Lemma follows the same procedure as Lemma 4.6 except the warped $\mu$-bubble has prescribed boundary. Refer the detailed proof of Theorem 4.4 in [Hong24] in ambient dimension $4$.
@@ 4.3. Almost linear volume growth
We show the almost linear growth.
Lemma 4.8 (Almost linear volume growth of a boundary or interior end).
Let $(X^{5},g)$ be a complete $5$-dimensional Riemannian manifold with bounded geometry and let $M^{4}\hookrightarrow(X^{5},g)$ be a simply-connected (strongly) stable CMC immersion (two-sided if it is minimal). Suppose $\operatorname{biRic}_{X}\geqslant 0$ and
$\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\varepsilon_{0}$
for a positive constant $\varepsilon_{0}$ and some $k\in[\frac{3}{2},2]$, where $\lambda_{k}$ is the infimum of the $k$-triRicci curvature. Let $\{E_{l}\}_{l\in\mathbb{N}}$ be an end of $M$ adapted to $p\in M$ with length scale $L$, where $L$ is the constant from Lemma 4.6 and Lemma 4.7. Let $M_{l}:=E_{l}\cap B_{(l+1)L}(p)$. Then there is a universal constant $C>0$ such that
$\operatorname{Vol}_{M}(M_{l})<C$
@@
The proof is very similar to that of [chodosh-li-stryker, Lemma 5.4]. For completeness, we sketch it here.
We apply Lemma 4.6 to the end $E_{l}$, which supplies an open set $\Omega_{l}$ in $B_{L/2}(\partial E_{l})\cap E_{l}$ with $\partial E_{l}\subset\partial\Omega_{l}$. Note that $\Omega_{l}\subset M_{l}$ and $\partial E_{l+1}\cap\overline{\Omega_{l}}=\emptyset$ by construction. Denote $\Sigma^{(i)}_{l}$ the components of $\Sigma_{l}:=\partial\Omega_{l}\backslash\partial E_{l}$.
As claimed in [chodosh-li-stryker, Lemma 5.4], $\overline{M_{l}}$ is connected and $\Sigma^{(i)}_{l}$ separates $\partial E_{l}$ and $\partial E_{l+1}$. By Lemma 4.6, we have $\operatorname{diam}(\Sigma_{l})\leqslant c$.
We claim that there is a constant $D>0$ independent of $l$ such that $\operatorname{diam}(M_{l})<D$. Let $q$ and $r$ be any points in $M_{l}$. The radial geodesic connecting them intersects $\partial E_{l-1}$ and $\partial E_{l}$. By the choice of the components of $\{\Sigma_{l}\}$, an arc of this curve of length at most $2L$ joins $q$ to $\Sigma_{l-1}$. The same argument applies to $r$. Therefore, the diameter of $M_{l}$ is bounded by $D:=4L+c$.
By Lemma 2.4, we know that $M$ has bounded second fundamental form (e.g. consider the comapct immersion of $B_{M}(q,1)$ for any $q\in M$). By Lemma 2.3, it implies that $\operatorname{Ric}_{M}$ is bounded from below. Hence, by Bishop–Gromov volume comparsion, there is a constant $C>0$ so that
$\operatorname{Vol}(B_{M}(x,D))\leqslant C$
for all $x\in M$. Since $M_{l}\subset\operatorname{Vol}(B_{M}(x,D))$ for any $x\in M_{l}$, the lemma follows. ∎

### v3

@@ 4.2. Volume bound for bubbles
$|\nabla h|\leqslant\varepsilon_{0}\left(1+\frac{\beta h^{2}}{\varepsilon_{0}}\right).$
We now consider the functional $\mathcal{A}_{k}$ on $M^{n}$, where the weight function $u$ satisfies
$\begin{split}-\Delta_{M}u&=\left(|A_{M}|^{2}+\operatorname{Ric}^{X}(\eta,\eta)\right)u~{}~{}\mbox{in}~{}~{}\mathring{M},\\ u&=1~{}~{}\mbox{on}~{}~{}\partial M;\end{split}$
see [Wu23, chen-hong-li-docarmoquestion] for more details on the existence of $u$. We assume that $\Omega$ is the minimizer of $\mathcal{A}_{k}$ and $\Sigma$ is a connected component of $\partial\Omega\backslash\partial M$. Let $\nu$ be the unit normal vector to $\Sigma$.
By Lemma 4.4, we have
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\geqslant&\int_{\Sigma}\left(|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-ku^{-1}\Delta_{M}u\right)u^{k}\psi^{2}\\ &+\int_{\Sigma}\nabla_{\nu}(u^{k}h)\psi^{2}-k(k-1)u^{k-2}(\nabla_{\nu}u)^{2}\psi^{2}.\end{split}$
Plugging $\phi=u^{\frac{k}{2}}\psi$ into the above inequality yields
$\int_{\Sigma}|\nabla\phi|^{2}+\frac{k^{2}}{4}u^{-2}$ $|\nabla u|^{2}\phi^{2}-ku^{-1}\phi\nabla\phi\cdot\nabla u-k\phi^{2}u^{-1}\Delta_{\Sigma}u$
$\geq$ $\int_{\Sigma}\left(|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-ku^{-1}\Delta_{M}u\right)\phi^{2}$
$+\int_{\Sigma}kh(\nabla_{\nu}\log u)\phi^{2}+\nabla_{\nu}h\phi^{2}-k(k-1)(\nabla_{\nu}\log u)^{2}\phi^{2}.$
Integration by parts yields
$\int_{\Sigma}|\nabla\phi|^{2}$ $+\left(\frac{k^{2}}{4}-k\right)u^{-2}|\nabla u|^{2}\phi^{2}+ku^{-1}\phi\nabla\phi\cdot\nabla u$
$\geq$ $\int_{\Sigma}\left(|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-ku^{-1}\Delta_{M}u\right)\phi^{2}$
$+\int_{\Sigma}kh(\nabla_{\nu}\log u)\phi^{2}+\nabla_{\nu}h\phi^{2}-k(k-1)(\nabla_{\nu}\log u)^{2}\phi^{2}.$
By Cauchy-Schwarz inequality, we obtain
$\frac{4}{4-k}\int_{\Sigma}|\nabla\phi|^{2}$ $\geq\int_{\Sigma}\left(|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-ku^{-1}\Delta_{M}u\right)\phi^{2}$
$+\int_{\Sigma}kh(\nabla_{\nu}\log u)\phi^{2}+\nabla_{\nu}h\phi^{2}-k(k-1)(\nabla_{\nu}\log u)^{2}\phi^{2}.$
Substituting $h$ by $H^{\Sigma}+k\nabla_{\nu}\log u$ yields
(4.8) $\frac{4}{4-k}\int_{\Sigma}|\nabla\phi|^{2}\geq$ $\int_{\Sigma}\left(|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-ku^{-1}\Delta_{M}u\right)\phi^{2}$
$+\int_{\Sigma}kH^{\Sigma}(\nabla_{\nu}\log u)\phi^{2}+\nabla_{\nu}h\phi^{2}+k(\nabla_{\nu}\log u)^{2}\phi^{2}.$
@@
(4.9) $-ku^{-1}\Delta_{M}u+|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu).$
We select a local orthonormal frame $\{e_{i}\}_{i=1}^{n-1}$ on $\Sigma^{n-1}$ and extend it by assigning $\nu=e_{n}$ and $\eta=e_{n+1}$ so that $\{e_{1},\cdots,e_{n-1},\nu,\eta\}$ is a local orthonormal frame of $X^{n+1}$. By Gauss equation, we have
$\begin{split}\operatorname{Ric}^{\Sigma}(e_{1},e_{1})&=\sum_{i=1}^{n-1}\operatorname{R}^{\Sigma}_{i1i1}=\sum_{i=1}^{n-1}\left(\operatorname{R}^{M}_{i1i1}+A^{\Sigma}_{ii}A^{\Sigma}_{11}-\left(A^{\Sigma}_{i1}\right)^{2}\right)\\ &=\sum_{i=1}^{n-1}\operatorname{R}^{M}_{i1i1}+H^{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{n-1}\left(A^{\Sigma}_{i1}\right)^{2}\\ &=\sum_{i=1}^{n-1}\operatorname{R}^{X}_{i1i1}+(H^{M}-A^{M}_{nn})A^{M}_{11}-\sum_{i=1}^{n-1}(A^{M}_{i1})^{2}+H^{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{n-1}\left(A^{\Sigma}_{i1}\right)^{2}.\end{split}$
On the other hand,
$\operatorname{Ric}^{M}(\nu,\nu)$ $=\sum_{i=1}^{n}\left(R^{X}_{inin}+A^{M}_{nn}A_{ii}^{M}-A_{ni}^{2}\right)$
$=\sum_{i=1}^{n}R^{X}_{inin}+H^{M}A^{M}_{nn}-\sum_{i=1}^{n}(A_{in}^{M})^{2}.$
@@
$\begin{split}-ku^{-1}&\Delta_{M}u+|A^{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)+\alpha\operatorname{Ric}^{\Sigma}(e_{1},e_{1})\\ =&k|A^{M}|^{2}+k\operatorname{Ric}^{X}(\eta,\eta)+\operatorname{Ric}^{M}(\nu,\nu)+|A^{\Sigma}|^{2}+\alpha\operatorname{Ric}^{\Sigma}(e_{1},e_{1})\\ =&k\sum_{i=1}^{n+1}\operatorname{R}^{X}_{i(n+1)i(n+1)}+\sum_{i=1}^{n}\operatorname{R}^{X}_{inin}+\alpha\sum_{i=1}^{n-1}\operatorname{R}^{X}_{i1i1}\\ &+k|A^{M}|^{2}+\alpha(H^{M}-A^{M}_{nn})A^{M}_{11}-\alpha\sum_{i=1}^{n-1}\left(A^{M}_{i1}\right)^{2}+H_{M}A^{M}_{nn}-\sum_{i=1}^{n}\left(A^{M}_{in}\right)^{2}\\ &+|A^{\Sigma}|^{2}+\alpha H^{\Sigma}A^{\Sigma}_{11}-\alpha\sum_{i=1}^{n-1}(A^{\Sigma}_{i1})^{2}\\ \geq&k\sum_{i=1}^{n+1}\operatorname{R}^{X}_{i(n+1)i(n+1)}+\sum_{i=1}^{n}\operatorname{R}^{X}_{inin}+\alpha\sum_{i=1}^{n-1}\operatorname{R}^{X}_{i1i1}\\ +&C(k,n)H_{M}^{2}+|A^{\Sigma}|^{2}+\alpha H^{\Sigma}A^{\Sigma}_{11}-\alpha\sum_{i=1}^{n-1}(A^{\Sigma}_{i1})^{2}\ \end{split}$
where the last inequality is due to Proposition 4.5 or Proposition 4.6 and $C(k,n)$ is defined in (4.7).
Now the inequality (4.8) becomes
$\frac{4}{4-k}$ $\int_{\Sigma}|\nabla\phi|^{2}\geq\int_{\Sigma}(-\alpha\operatorname{Ric}^{\Sigma}(e_{1},e_{1}))\phi^{2}$
$+\int_{\Sigma}\left(k\sum_{i=1}^{n+1}\operatorname{R}^{X}_{i(n+1)i(n+1)}+\sum_{i=1}^{n}\operatorname{R}^{X}_{inin}+\alpha\sum_{i=1}^{n-1}\operatorname{R}^{X}_{i1i1}+C(k,n)H_{M}^{2}\right)\phi^{2}$
$+\int_{\Sigma}\left(|A_{\Sigma}|^{2}+\alpha H^{\Sigma}A^{\Sigma}_{11}-\alpha\sum_{i=1}^{n-1}(A^{\Sigma}_{i1})^{2}\right)\phi^{2}$
$+\int_{\Sigma}kH^{\Sigma}(\nabla_{\nu}\log u)\phi^{2}+\nabla_{\nu}h\phi^{2}+k(\nabla_{\nu}\log u)^{2}\phi^{2}.$
We now estimate the following quantity
$I:=|A^{\Sigma}|^{2}+\alpha H^{\Sigma}A^{\Sigma}_{11}-\alpha\sum_{i=1}^{n-1}(A^{\Sigma}_{i1})^{2}+kH^{\Sigma}(\nabla_{\nu}\log u)+\nabla_{\nu}h+k(\nabla_{\nu}\log u)^{2}$
and expect that $I\geq\beta h^{2}$ for some $\beta>0$. In the same spirit in [mazet, Page 11-12], it is equivalent to show that the matrix
(4.10) $G=\begin{pmatrix}\frac{1}{n-1}+\frac{\alpha}{n-1}-\frac{\alpha}{(n-1)^{2}}+\frac{1}{k}-1&\frac{\alpha}{2}\sqrt{\frac{n-2}{n-1}}(1-\frac{2}{n-1})&\frac{1}{2}-\frac{1}{k}\\ \frac{\alpha}{2}\sqrt{\frac{n-2}{n-1}}(1-\frac{2}{n-1})&1-\alpha\frac{n-2}{n-1}&0\\ \frac{1}{2}-\frac{1}{k}&0&\frac{1}{k}-\beta\end{pmatrix}$
is positive semi-definite for some $k$, $\alpha$ and $\beta$. If so, by our assumptions, we obtain that
$\frac{4}{4-k}\int_{\Sigma}|\nabla\phi|^{2}$ $\geq\int_{\Sigma}(-\alpha\operatorname{Ric}^{\Sigma}(e_{1},e_{1})+2\varepsilon_{0}+\beta h^{2}+\nabla_{\nu}h)\phi^{2}.$
@@
According to the definition of $h$ and choosing $e_{1}$ realizing the minimum Ricci curvature $\lambda_{\operatorname{Ric}}(\Sigma)=\operatorname{Ric}_{\Sigma}(e_{1},e_{1})$, we get
@@
$\begin{split}\frac{4}{4-k}\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\geqslant\int_{\Sigma}\left(-\alpha\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)\phi^{2}.\end{split}$
@@
Therefore, there is a positive smooth function $\theta\in C^{\infty}(\Sigma)$ so that
@@
$\gamma\Delta_{\Sigma}\theta\leq\lambda_{\operatorname{Ric}}\theta-\frac{\varepsilon_{0}}{\alpha}\theta,\quad\gamma=\frac{4}{(4-k)\alpha}.$
We need
$\gamma\leq\frac{n-2}{n-3}$
to hold so that we can apply [antonelli-xu, Theorem 1] to conclude that there exists a constant $c$ such that $\operatorname{Vol}(\Sigma)\leq c.$
@@ 4.3. Almost linear volume growth
In this part, we will upgrade the volume estimate of $\mu$-bubbles to the volume growth of ends in stable CMC hypersurfaces under several curvature conditions.
Theorem 4.9.
Let $(X^{n+1},g)$ $(4\leq n\leq 5)$ be a complete $(n+1)$-dimensional Riemannian manifold, and let $M^{n}\hookrightarrow(X^{n+1},g)$ be a simply-connected stable two-sided CMC hypersurface. Let $\{E_{l}\}_{l\in\mathbb{N}}$ be an end of $M^{n}$ adapted to $p\in M$ with length scale $L$, where $L$ is the constant from Lemma 4.7. Denote $M_{l}:=E_{l}\cap B_{(l+1)L}(p)$. Assume that either $X$ has bounded geometry or $X$ has weakly bounded geometry and $\operatorname{Ric}^{X}_{n-1}\geq C_{2}$ for a constant $C_{2}$. Under assumptions in Lemma 4.7, there is a universal constant $C>0$ independent of $l$ such that
$\operatorname{Vol}(M_{l})<C$
@@
We apply Lemma 4.7 to the end $E_{l}$, which supplies an open set $\Omega_{l}$ in $B_{L/2}(\partial E_{l})\cap E_{l}$ with $\partial E_{l}\subset\partial\Omega_{l}$. Note that $\Omega_{l}\subset M_{l}$ and $\partial E_{l+1}\cap\overline{\Omega_{l}}=\emptyset$ by construction.
As claimed in [chodosh-li-stryker, Lemma 5.4], $\overline{M_{l}}$ is connected and there exists a component $\Sigma^{*}_{l}$ of the $\mu$-bubble $\Sigma_{l}:=\partial\Omega_{l}\backslash\partial E_{l}$ separating $\partial E_{l}$ and $\partial E_{l+1}$. In fact, if not, we can find a non-contractible closed curve, contradicting the simply-connectedness of $M^{n}$. By Lemma 4.7, we have $\operatorname{Vol}(\Sigma_{l}^{*})\leqslant c$ where $c$ is independent of $l.$
We now claim that there is a constant $D>0$ independent of $l$ such that $\operatorname{Vol}(M_{l})<D$. It can be deduced by using the area comparison theorem of hypersurfaces. This idea was firstly applied by Wei-Xu-Zhang in [Wei_Xu_Zhang]. Due to the fact that $\Sigma_{l-1}^{*}$ separates $p$ and $E_{l}$, every hypersurface $\partial B_{s}(p)\cap M_{l}$ for $s\in[lL,(l+1)L]$ lies behind $\Sigma_{l-1}^{*}$, which means any geodesic starting from $p$ to $\partial B_{s}(p)\cap M_{l}$ goes through $\Sigma_{l-1}^{*}$.
It follows from Lemma 2.5 and Lemma 2.3 that $\operatorname{Ric}^{M}\geq-Cg$. Then by Bishop–Gromov volume comparison theorem, we have
$\operatorname{Vol}(\partial B_{s}(p)\cap M_{l})\leq C\frac{e^{(n-1)s}}{e^{(n-1)(l-1)L}}\operatorname{Vol}(\Sigma_{l-1}^{*})\leq Ce^{2(n-1)L}$
for any $s\in[lL,(l+1)L]$. Integrating and using co-area formula yield
$\operatorname{Vol}(M_{l})\leq CLe^{2(n-1)L}.$